8) Calculate the yield to maturity for the following bond quote 7 1/16 28 Close 1095/8.
please show calculations

that the entire question
Answer #1

Yield to Maturity

= C + FV - PV /n/FV+PV/2

= 7+ 109-28/0.67/109+28/2

= 7 + 81/0.67 / 137/2

=7+120.90 / 68.5

= 127.90 / 68.5

= 1.87 %